# Distinctive Feature Codec: Adaptive Segmentation for Efficient Speech Representation

### v1

@@ 4 Experimental Results
Speech Tokenizer Pre-train [18] RVQ 50 50 500 0.9505 2.4454 1.3921 0.6412 0.0755
RVQ 50 100 1000 0.4011 1.1522 1.9159 0.7739 0.0680
RVQ 50 400 4000 0.1139 0.3223 3.1594 0.9196 0.0442
Speech Tokenizer RVQ 10 10 100 0.4487 2.0183 1.2844 0.6695 0.9340
RVQ 12.5 12.5 125 0.4290 1.9448 1.3245 0.6624 0.7701
RVQ 20 20 200 0.3796 1.7401 1.4960 0.7225 0.6887
RVQ 50 50 500 0.2342 0.5648 2.4496 0.8439 0.1698
FSQ 20 20 320 0.1933 0.4699 2.5891 0.8615 0.1428
Distinctive Codec RVQ 9.5 9.5 95 0.4481 2.0040 1.3312 0.6930 0.8523
FSQ 9.5 9.5 152 0.4033 1.9042 1.4649 0.7049 0.6794
GSQ 9.5 9.5 152 0.2857 1.3213 1.9147 0.7675 0.4265
GSQ 15.7 15.7 251 0.2468 0.9072 2.3092 0.8203 0.2637
We explore the capabilities and properties of Distinctive Features through experiments with our base codec model. Our model shown in Table 1produces tokens at an average rate of 9.5 Hz and 15.7 Hz, significantly lower than standard frame-based approaches.
Table 1 presents a comparison between our Distinctive Codec and Speech Tokenizer across various configurations. For the Speech Tokenizer Pre-train results, we directly used their released model for inference, extracting results using the first RVQ layer, second RVQ layer, and all eight RVQ layers respectively. Importantly, our implementation of Speech Tokenizer does not include HuBERT distillation to ensure a fair architectural comparison.
The effectiveness of our Distinctive Features approach is evident when comparing the RVQ results between our models. Specifically, our Distinctive Codec operating at 9.5 Hz outperforms Speech Tokenizer’s 10 Hz configuration across all key metrics. Despite using a slightly lower token rate, our approach reduces MEL Error by 0.0006, decreases STFT distortion by 0.0143, improves perceptual quality with PESQ by 0.0468, enhances intelligibility with STOI by 0.0235, and reduces WER by 0.0817. These consistent gains demonstrate that dynamic token allocation based on distinctive features provides clear advantages over the uniform frame-based processing employed in Speech Tokenizer.
Further improvements are observed when replacing RVQ with FSQ in our Distinctive Codec. However, we found that Speech Tokenizer becomes unstable with FSQ at lower token rates, preventing a direct comparison at our model’s operating point of 9.5 Hz.
The comparison between GSQ and FSQ at the same token rate (9.5 Hz) clearly demonstrates the effectiveness of our Group-wise Scalar Quantization approach. GSQ substantially outperforms FSQ across all metrics, reducing MEL error by 0.1176, decreasing STFT distortion by 0.5829, improving PESQ by 0.4498, increasing STOI by 0.0626, and lowering WER by 0.2529. This validates our hypothesis that decomposing high-dimensional quantization into multiple lower-dimensional sub-problems enhances stability and representation quality for variable-length segments. Additionally, in Appendix H, our model demonstrates strong generalization on unseen languages, and Appendix I confirms our approach outperforms other SOTA models at comparable token rates.
5 Ablation Study and Discussion
5.1 Analysis of Distinctive Feature Effectiveness
Theoretical Derivation Hypothesis
where $g$ is the decoder and $Q$ is the quantizer that maps $z$ to a finite codebook. High-resolution quantization theory [63] approximates this distortion as:
[…]
Codebook Analysis
Table 2: Comparison between FSQ and our proposed GSQ(many to many) at different reconstruction levels. Both quantization methods are evaluated on the same Distinctive Codec architecture operating at 9.5 Hz.
[…]
FSQ 32 0.1768 0.5844 2.5467 0.8645 0.1301
FSQ 64 0.1407 0.4330 2.7909 0.8950 0.0732
GSQ 24 0.2056 0.6963 2.3578 0.8474 0.1794
GSQ 32 0.1872 0.6207 2.4732 0.8597 0.1433
GSQ 64 0.1456 0.4423 2.7848 0.8889 0.0749
5.2 Analysis of Group-wise Scalar Quantization Effectiveness
Comparative Analysis of Quantization Methods
Building on our theoretical analysis of distinctive features, we next investigate the effectiveness of our quantization strategy when applied to segments of varying complexity. Table 2 compares our GSQ approach with standard FSQ across different reconstruction levels. While our primary results focused on compressing variable-length segments into single tokens, this analysis examines how effectively both methods preserve information at different dimensionalities. The reconstruction level represents the latent space dimension to which each segment is compressed before quantization. Our GSQ approach offers a significant advantage by dramatically reducing the effective dictionary size required for high-quality representations. Whereas direct application of FSQ to high-dimensional features would demand an extremely large dictionary, our group-wise strategy effectively partitions the feature space into multiple specialized quantization subproblems. This approach not only enhances computational efficiency but also improves representation stability for distinctive feature-based segments. Despite using a much smaller effective dictionary, GSQ achieves performance comparable to FSQ across all metrics, particularly at higher reconstruction levels.
(a) FSQ latent space clustering
(b) GSQ latent space clustering
Figure 5: t-SNE visualization [64] of latent space clusters. (a) Standard FSQ shows overlapping clusters. (b) GSQ produces more distinct clusters with clearer boundaries.
Analysis of GSQ’s Effectiveness Through Latent Space Geometry
Our information-theoretic analysis suggests that if GSQ preserves more structural information, this should be reflected in the geometric organization of the latent space [65, 66]. Specifically, a quantization method that maintains higher mutual information and lower distortion should produce more coherent and well-separated clusters when visualized. To test this hypothesis, we employed t-SNE visualization [64], which maps high-dimensional data to two dimensions while preserving local neighborhood relationships, making it ideal for assessing how well structural information is maintained after quantization. Figure 5 shows striking differences between standard FSQ (Figure 5(a)) and our GSQ approach (Figure 5(b)). GSQ produces significantly more coherent clusters with clearer boundaries, indicating that GSQ preserves class structure and separability in the latent space. These results directly validate our hypothesis: GSQ’s decomposition strategy prevents information mixing across feature groups, enabling it to preserves more meaningful information within severe token constraints. The improved geometric organization of GSQ’s latent space explains its superior reconstruction quality across all our evaluation metrics. The details of experiments in the Appendix D
6 Conclusion
In this paper, we investigated the effectiveness of distinctive features for speech representation, demonstrating that identifying and prioritizing perceptually significant transitions leads to more efficient tokenization than conventional frame-based processing. Our experiments address the fundamental limitation of uniform processing approaches that ignore the non-uniform information distribution in speech, showing that distinctive feature-based segmentation produces more coherent latent representations with improved codebook utilization. The theoretical and empirical advantages of aligning segment boundaries with natural acoustic transitions validate principles from linguistic theory while extending them into modern deep learning frameworks. Our proposed GSQ strategy provides stable quantization at low frame rates through decomposing high-dimensional features into specialized subproblems. This work establishes distinctive features as a promising direction for neural speech representation that could benefit large-scale speech language models and potentially extend to other continuous signal domains.
Limitation
Performance characteristics may vary when implementing our approach on different hardware platforms, as our experiments were conducted exclusively using NVIDIA V100 GPUs as the computational infrastructure. While we have provided detailed model configurations, hyperparameters, and training procedures to ensure reproducibility, future work could explore the scalability and efficiency of our distinctive feature-based codec across various computational environments, including more recent GPU architectures or specialized audio processing hardware.
Appendix A Analysis of Different Types of Speech Codecs
A.1 Case Study of Sylber Codec
Table 3: Analysis of Sylber Codec performance. The table shows mel_error, stft, pseq, and stoi metrics for different KM (K-Means) cluster sizes.
Sylber Codec mel_error $\downarrow$ stft $\downarrow$ pseq $\uparrow$ stoi $\uparrow$
KM=5K 0.6204 2.0246 1.0025 0.7066
KM=10K 0.6155 2.2049 1.2501 0.7136
KM=20K 0.6727 2.1525 0.8135 0.7003
Let us examine Sylber [47], a widely recognized flow matching-based codec, as a case study. In the original work, Sylber demonstrated the ability to compress the token rate to remarkably low levels while exhibiting improved model capabilities with increasing K-means cluster sizes. The authors reported significant improvements in recognition metrics like WER as the vocabulary size expanded.
However, despite these impressive compression capabilities, subsequent research has shown that Sylber’s representations do not perform ideally when used to train large language models. We hypothesize that this limitation stems from the excessive strength of the generative model, which results in discrete representations that lack sufficient original speech information. The powerful generative capabilities of flow matching-based models essentially "guess" or hallucinate speech content rather than preserving it in the tokens.
This phenomenon is evident in Table 3, where we observe that as the model’s capacity increases from KM=10K to KM=20K, the perceptual speech quality score (pseq) actually decreases from 1.2501 to 0.8135. This counterintuitive result suggests that while the model may generate high-quality speech, this speech is not necessarily similar to the original audio input. The model appears to be relying more on its generative capabilities than on preserving the information contained in the original speech signal.
Given these considerations, we focus our experimental investigation primarily on end-to-end trained codecs, which we believe provide a more direct and transparent approach to learning speech representations with higher information preservation.
A.2 Empirical Analysis to Determine Optimal Token Rate
The quest for determining an optimal token rate in speech representation presents a fundamental challenge in codec design. While models like SpeechTokenizer operate at relatively high rates (50 Hz) and syllabic approaches like Sylber achieve much lower rates (around 5 Hz), establishing a theoretically sound minimum rate remains unresolved. To address this challenge, we can leverage insights from text tokenization, where linguistic content typically requires 5-6 tokens per second. However, speech contains additional information beyond text, including prosody, speaker characteristics, and emotional cues. Determining how much this extra information increases the required token rate necessitates measuring the information content in speech signals. Unfortunately, directly measuring information content in continuous speech signals is inherently difficult [67], requiring alternative approaches to estimate the necessary token rate for complete speech representation. To address this challenge, we developed an indirect method for estimating speech information density by analyzing the relationship between different types of representation vectors derived from speech. This approach allows us to circumvent the difficulties of direct information measurement while still obtaining meaningful insights about optimal token rates.
Construction of Specialized Representation Vectors
We constructed two distinctive representation vectors to capture different aspects of speech information:
1. A speech-only representation vector designed to maximize similarity with HuBERT [27] features:
$\mathcal{J}_{\text{speech}}=-\frac{\mathbf{r}\cdot\mathbf{h}_{\text{ssl}}}{\|\mathbf{r}\|\|\mathbf{h}_{\text{ssl}}\|}$ (15)
where $\mathbf{h}_{\text{ssl}}$ represents the weighted sum of HuBERT features across layers.
2. A text-filtered representation vector constructed through a contrastive learning objective:
$\mathcal{J}_{\text{filtered}}=-\frac{1}{B}\sum_{i=1}^{B}\log\frac{\exp(\text{sim}(\mathbf{r},\mathbf{h}_{\text{ssl}}^{(i)})/t)}{\exp(\text{sim}(\mathbf{r},\mathbf{h}_{\text{ssl}}^{(i)})/t)+\exp(\text{sim}(\mathbf{r},\mathbf{s}_{\text{text}}^{(i)})/t)}$ (16)
Where $\mathbf{s}_{\text{text}}$ represents sentence embeddings from transcriptions and $t$ is a temperature parameter.
Information Content Estimation
To quantify the relative information content in these representations, we employed a neural information estimation approach [68] to measure mutual information between our representation vectors and sentence embeddings:
$I_{\Omega}(\mathbf{r},\mathbf{s})=\sup_{\omega\in\Omega}\mathbb{E}_{P(\mathbf{r},\mathbf{s})}[F_{\omega}(\mathbf{r},\mathbf{s})]-\log\mathbb{E}_{P(\mathbf{r})P(\mathbf{s})}[e^{F_{\omega}(\mathbf{r},\mathbf{s})}]$ (17)
This estimation technique enables us to compare the linguistic information preserved in speech-only versus text-filtered representations, providing insight into the information density of speech signals beyond what is captured in text alone.
Results and Implications for Token Rate
Our analysis revealed a significant finding: the mutual information between speech-only representations and sentence embeddings was approximately 1.4 times higher than between text-filtered representations and the same embeddings. This ratio provides a quantifiable measure of the additional information present in speech beyond linguistic content.
Given that English text typically requires 5-6 tokens per second for adequate representation, and syllabic models like Sylber operate within this range, our information ratio suggests that a more complete speech representation would require approximately 7-8.4 tokens per second (applying our 1.4× multiplier) to adequately capture the full information content present in speech signals. Notably, our empirical findings align with this theoretical estimate, as we observed that our model becomes unstable and fails to train effectively when the token rate falls below 9 tokens per second.
It’s important to note that this experiment is heuristic in nature, designed to provide insight rather than precise measurements. Nevertheless, these findings help explain our observations with flow-matching models operating at very low token rates. While these models often produce intelligible speech, they tend to deviate noticeably from the original audio input in terms of overall acoustic fidelity. Our analysis suggests that when the token rate falls significantly below our estimated threshold, the generative model is effectively "hallucinating" aspects of speech rather than encoding them in the tokens, as there is insufficient capacity to preserve the full information content of the original signal.
Through this information-theoretic approach, we establish a principled foundation for token rate selection in speech codec design, moving beyond purely empirical determination of compression rates and providing theoretical insights into optimal token rate selection, even if the exact values may require further validation.
Appendix B Details of Model Configurations
B.1 Details of Distinctive Feature Detector
@@
The detector uses a 5-layer CNN structure processing raw audio input directly. Each layer consists of a 1D convolutional operation followed by batch normalization and LeakyReLU activation. The network employs variable kernel sizes and strides to progressively downsample the input while capturing acoustic patterns at different time scales. For our primary configuration yielding 9.5 tokens per second (as shown in Table 1), the convolutional layers use kernel sizes of 10, 8, 8, 4, and 4, with corresponding stride values of 5, 4, 4, 2, and 2. For our higher frame rate configuration (15.7 tokens per second), we adjusted the stride values to 5, 4, 4, 4, and 2, demonstrating the adaptability of our approach. This flexibility allows our distinctive feature-based method to operate effectively across different token rates, unlike approaches such as Sylber Codec that are limited to specific syllable-level rates. Our configurable architecture provides an effective receptive field capable of capturing both local and broader acoustic transitions while allowing token rate adjustments based on application requirements.
@@
Boundary Detection Inference Algorithm
@@
This boundary detection process enables our codec to adaptively identify regions where speech characteristics undergo significant changes, rather than relying on fixed-interval processing. The detected boundaries are then used to guide the variable-length segmentation in our Distinctive Codec framework.
B.2 Details of Distinctive Codec
@@
Boundary Detection Integration
During inference, the boundary detector identifies perceptually significant transitions in the input speech. These boundaries are then used to segment the encoded representations before compression. For each input waveform, the boundary detection process works as follows:
The detector processes the raw audio input and outputs similarity scores that undergo normalization. Boundary detection then identifies local maxima in the processed similarity scores using prominence, width, and distance parameters that were automatically tuned during training. The detected boundaries guide the variable-length segmentation in our Distinctive Codec framework.
For each segment defined by these boundaries, the DFE encodes the variable-length features into a fixed-length representation. These compressed representations are then quantized using our GSQ approach. During decoding, the DFD expands each quantized representation back to its original length based on the segment information stored during encoding. Finally, the reconstructed segments are concatenated and processed through the decoder to generate the output waveform.
@@
Appendix C Details of Evaluation Metrics
We evaluate our Distinctive Codec across three key dimensions:
Reconstruction Quality
We employ multiple complementary metrics to assess how faithfully the codec reproduces the original speech signal:
  • Mel-spectral Error: Measures the L1 or L2 distance between the original and reconstructed speech in the mel-spectral domain, correlating well with perceptual differences.
  • STFT Distance: Evaluates the accuracy of the time-frequency representation, capturing both phase and magnitude discrepancies across multiple frequency bands.
  • PESQ: The Perceptual Evaluation of Speech Quality provides a standardized objective assessment of perceived audio quality on a scale from -0.5 to 4.5, with higher scores indicating better quality.
Intelligibility
We quantify how well the encoded-decoded speech preserves linguistic content:
  • Word Error Rate (WER): Measured using the Whisper en-medium model, following the evaluation protocol in SpeechTokenizer [18]. Lower WER indicates better preservation of intelligible speech content, which is critical for downstream speech applications.
  • STOI: Short-Time Objective Intelligibility measures the correlation between the temporal envelopes of the original and processed speech signals. Values range from 0 to 1, with higher values indicating better intelligibility. This metric is particularly sensitive to distortions that affect speech understanding.
Encoding Efficiency
We track the compression performance through:
  • Token Ratio (TKR): Indicates the number of tokens required to represent one second of speech at 16 kHz sampling rate. Lower values indicate more efficient representation. For example, if frame rate is 50hz and residual of RVQ is 8, the TKR should be 400.
  • Bits Per Second (BPS): Quantifies the bitrate of the compressed speech. Following the approach in [55], we calculate BPS as:
    $\text{BPS}=f_{r}\sum_{i=0}^{n}\lceil\log_{2}(k_{i})\rceil$ (18)
    where $f_{r}$ is the frame rate, $n$ is the number of Group-FSQs or residual level, and $k_{i}$ is the codebook size.
Appendix D Details of Ablation Study and Analysis
For our cluster analysis experiment comparing FSQ and GSQ representations, we collected frame-level features from our test set and processed them through both quantization methods. We randomly sampled 500 frames from the quantized output representations to ensure computational efficiency while maintaining statistical significance.
We performed K-means clustering with K=4 on the sampled features using a maximum of 30 iterations and a fixed random seed (42) for reproducibility. To evaluate cluster quality, we calculated the silhouette score, which measures the cohesion within clusters and separation between clusters on a scale from -1 to 1.
For visualization purposes, we applied t-SNE dimensionality reduction with perplexity=30 to project the high-dimensional features onto a 2D plane. The resulting visualizations used distinct colors (red, green, blue, and gold) with alpha transparency of 0.6 to better reveal density patterns. We also analyzed cluster size distributions using histograms to identify potential quantization biases or degenerate clustering solutions.
D.1 Details of Semantic Information Distillation

### v2

@@ III-C Group-wise Scalar Quantization
14:return $\hat{x}$
IV Experimental Setup
We implement our Distinctive Codec using the widely-adopted SEANet-based [tagliasacchi2020seanet] encoder-decoder architecture, which has become the de facto standard in modern neural speech codecs including SoundStream [zeghidour2021soundstream], EnCodec [kumar2023high], DAC [kumar2023high], and SpeechTokenizer [zhang2024speechtokenizer]. This architectural choice enables direct comparison with existing methods while isolating the impact of our distinctive feature-based adaptive segmentation versus conventional fixed-rate processing. The encoder employs strided convolutional layers for downsampling, bidirectional LSTM layers for temporal modeling, and residual blocks for feature refinement. The decoder mirrors this structure to reconstruct the waveform from the latent representation.
For training and evaluation, we use the LibriSpeech dataset [panayotov2015librispeech], a widely-adopted benchmark containing 960 hours of read English speech with diverse speakers and acoustic conditions. We train on the standard train set and evaluate codec reconstruction quality on 500 randomly selected samples from the test-clean set. This evaluation protocol maintains consistency with prior codec research and enables direct performance comparison across methods. For downstream evaluation of depression detection, we use the DAIC-WOZ dataset [devault2014simsensei], which contains clinical interviews specifically designed for mental health assessment and provides naturalistic speech with diagnostically relevant temporal dynamics.
IV-A Details of Distinctive Feature Detector
@@
The detector uses a 5-layer CNN structure processing raw audio input directly. Each layer consists of a 1D convolutional operation followed by batch normalization and LeakyReLU activation. The network employs variable kernel sizes and strides to progressively downsample the input while capturing acoustic patterns at different time scales. For our primary configuration which yields 9.5 tokens per second (as shown in Table II), the convolutional layers use kernel sizes of 10, 8, 8, 4, and 4, with corresponding stride values of 5, 4, 4, 2, and 2. For our higher frame rate configuration (15.7 tokens per second), we adjusted the stride values to 5, 4, 4, 4, and 2, demonstrating the adaptability of our approach. This flexibility allows our distinctive feature-based method to operate effectively across different token rates, unlike approaches such as Sylber Codec that are limited to specific syllable-level rates. Our configurable architecture provides an effective receptive field capable of capturing both local and broader acoustic transitions while allowing token rate adjustments based on application requirements.
@@
Boundary Detection Inference
@@
IV-B Details of Distinctive Codec
@@
TABLE II: Performance comparison of frame-based and distinctive feature-based speech tokenization methods. Frame: frame rate (Hz); TKR: tokens per second [du2024funcodec]; BPS: bits per second [parker2024scaling]. RVQ: Residual Vector Quantizer [van2017neural]; FSQ: Finite Scalar Quantization [mentzerfinite]; GSQ: Group-wise Scalar Quantization (ours). Lower values are better for MEL Error, STFT, and WER; higher values are better for PESQ [pesq] and STOI [jensen2016algorithm]. All models use identical SEANet-based encoder-decoder architecture trained on LibriSpeech.
Segmentation Quantization Frame TKR BPS Metrics
[…]
Frame-based RVQ 10 10 100 0.4487 2.0183 1.2844 0.6695 0.9340
RVQ 12.5 12.5 125 0.4290 1.9448 1.3245 0.6624 0.7701
RVQ 20 20 200 0.3796 1.7401 1.4960 0.7225 0.6887
RVQ 50 50 500 0.2342 0.5648 2.4496 0.8439 0.1698
FSQ 20 20 320 0.1933 0.4699 2.5891 0.8615 0.1428
Distinctive Feature -based (Ours) RVQ 9.5 9.5 95 0.4481 2.0040 1.3312 0.6930 0.8523
FSQ 9.5 9.5 152 0.4033 1.9042 1.4649 0.7049 0.6794
GSQ 9.5 9.5 152 0.2857 1.3213 1.9147 0.7675 0.4265
GSQ 15.7 15.7 251 0.2468 0.9072 2.3092 0.8203 0.2637
@@
IV-C Codec Evaluation Metrics
We evaluate our Distinctive Codec using metrics for reconstruction quality, intelligibility, and encoding efficiency following previous works [zeghidour2021soundstream, defossezhigh]. For reconstruction, we measure mel-spectral error, STFT distance, and PESQ (Perceptual Evaluation of Speech Quality) scores [pesq]. For intelligibility, we measure Word Error Rate (WER) using the Whisper en-medium model [radford2023robust], following SpeechTokenizer [zhang2024speechtokenizer], and STOI (Short-Time Objective Intelligibility) [jensen2016algorithm] to quantify how accurately the speech content is preserved. We also track encoding efficiency through Token Ratio (TKR) [du2024funcodec], representing tokens per second of 16 kHz audio, and Bits Per Second (BPS). The BPS calculation follows the approach in [parker2024scaling] which will consider the vocabulary size of quantization method.
IV-D Depression Detection Evaluation
Evaluating the downstream performance of speech tokenizers for clinical applications presents fundamental methodological challenges. Training full-scale clinical assessment models requires substantial computational resources and carefully curated clinical data, making comprehensive comparison across multiple tokenization approaches prohibitively expensive. Furthermore, traditional codec evaluation metrics that focus on reconstruction quality and speech recognition fail to capture the preservation of subtle acoustic and temporal characteristics essential for mental health assessment, creating a significant evaluation gap between codec performance and clinical utility.
To enable systematic comparison of how different tokenization strategies preserve clinically relevant temporal information, we adopt the token projection evaluation framework proposed in [deshmukh2023pengi]. This methodology isolates the effects of tokenization by maintaining identical downstream architectures and training procedures across all evaluated methods, ensuring that performance differences directly reflect the quality of information preserved in discrete representations rather than variations in model design. We apply this comparative framework to the binary depression classification task using the DAIC-WOZ dataset [devault2014simsensei], which requires capturing subtle prosodic variations, pause patterns, and speech rate fluctuations—the temporal dynamics that our distinctive feature-based approach is designed to preserve.
Evaluation Protocol.
[…]
V Experimental Results
V-A Codec Results
Table II presents a systematic comparison between frame-based and distinctive feature-based tokenization approaches under identical architectural and training conditions. Our distinctive feature-based codec operates at average token rates of 9.5 Hz and 15.7 Hz, significantly lower than conventional fixed-rate processing while maintaining superior reconstruction quality.
The effectiveness of our distinctive feature-based approach is evident when comparing models at similar token rates. At 9.5 Hz, our method with RVQ quantization outperforms the frame-based baseline at 10 Hz across all metrics. Despite using a slightly lower token rate, our approach reduces MEL Error by 0.0006, decreases STFT distortion by 0.0143, improves perceptual quality (PESQ) by 0.0468, enhances intelligibility (STOI) by 0.0235, and reduces WER by 0.0817. These consistent improvements demonstrate that adaptive segmentation guided by acoustic boundaries preserves more acoustic information than arbitrary fixed-interval processing, even when using fewer tokens per second.
We observe further improvements when replacing RVQ with FSQ in our distinctive feature-based codec. Notably, frame-based approaches exhibit instability with FSQ at lower token rates (below 20 Hz), preventing direct comparison at our model’s operating point of 9.5 Hz. This instability supports our hypothesis that fixed-rate segmentation at long intervals forces the quantizer to represent acoustically heterogeneous content within single frames, leading to training difficulties. The comparison between GSQ and FSQ at the same token rate (9.5 Hz) clearly demonstrates the effectiveness of our Group-wise Scalar Quantization approach. GSQ substantially outperforms FSQ across all metrics, reducing MEL error by 0.1176, decreasing STFT distortion by 0.5829, improving PESQ by 0.4498, increasing STOI by 0.0626, and lowering WER by 0.2529. This validates our hypothesis that decomposing high-dimensional quantization into multiple lower-dimensional sub-problems enhances stability and representation quality for variable-length segments generated by distinctive feature-based segmentation.
V-B Depression Detection Results
TABLE III: Depression detection performance comparison on DAIC-WOZ development set. All models use identical downstream architecture (Llama 3.1 8B with projection module) and training protocol. Results are averaged over three random seeds.
Tokenization Method F1-Score
Frame-based (10Hz RVQ) 0.471
Distinctive Feature-based (9.5Hz RVQ) 0.533
Distinctive Feature-based (9.5Hz GSQ) 0.636
Table III presents the depression detection performance comparison across tokenization approaches. Our distinctive feature-based codec with GSQ achieves an F1-score of 0.636, substantially outperforming both the frame-based baseline (F1=0.471, +35.0% relative improvement) and our method with standard RVQ quantization (F1=0.533, +19.3% relative improvement). These results provide direct evidence that timing-preserving tokenization captures clinically relevant temporal dynamics that are destroyed by fixed-rate processing.
The performance hierarchy reveals two complementary effects. First, comparing frame-based (F1=0.471) versus distinctive feature-based with RVQ (F1=0.533) isolates the contribution of adaptive segmentation: aligning token boundaries with acoustic transitions preserves the pause patterns, speech rate variations, and prosodic contours that serve as established biomarkers for depression. Second, the further gain from GSQ (F1=0.636) demonstrates that stable quantization of variable-length segments is crucial for preserving fine-grained temporal information at low token rates. Together, these findings validate that our approach—adaptive segmentation plus robust quantization—effectively encodes the temporal dynamics essential for clinical speech analysis.
VI Ablation Study and Discussion
VI-A Analysis of Distinctive Feature Effectiveness
Theoretical Derivation Hypothesis
where $g$ is the decoder and $Q$ is the quantizer that maps $z$ to a finite codebook. High-resolution quantization theory [gray1998quantization] approximates this distortion as:
$D\approx G(d)\left(\int p(z)^{\frac{d}{d+2}}dz\right)^{\frac{d+2}{d}}2^{-\frac{2R}{d}}$ (14)
Codebook Analysis
VI-B Analysis of Group-wise Scalar Quantization Effectiveness
TABLE IV: Comparison between FSQ and our proposed GSQ (many to many) at different reconstruction levels. Both methods are evaluated on the same Distinctive Codec architecture operating at 9.5 Hz.
[…]
FSQ 32 0.1768 0.5844 2.5467 0.8645 0.1301
FSQ 64 0.1407 0.4330 2.7909 0.8950 0.0732
GSQ 24 0.2056 0.6963 2.3578 0.8474 0.1794
GSQ 32 0.1872 0.6207 2.4732 0.8597 0.1433
GSQ 64 0.1456 0.4423 2.7848 0.8889 0.0749
(a) FSQ latent space clustering
(b) GSQ latent space clustering
Figure 5: t-SNE visualization [van2008visualizing] of latent space clusters. (a) Standard FSQ shows overlapping clusters. (b) GSQ produces more distinct clusters with clearer boundaries.
Comparative Analysis of Quantization Methods
Building on our theoretical analysis of distinctive features, we next investigate the effectiveness of our quantization strategy when applied to segments of varying complexity. Table IV compares our GSQ approach with standard FSQ across different reconstruction levels. While our primary results focused on compressing variable-length segments into single tokens, this analysis examines how effectively both methods preserve information at different dimensionalities. The reconstruction level represents the latent space dimension to which each segment is compressed before quantization. Our GSQ approach offers a significant advantage by dramatically reducing the effective dictionary size required for high-quality representations. Whereas direct application of FSQ to high-dimensional features would demand an extremely large dictionary, our group-wise strategy effectively partitions the feature space into multiple specialized quantization subproblems. This approach not only enhances computational efficiency but also improves representation stability for distinctive feature-based segments. Despite using a much smaller effective dictionary, GSQ achieves performance comparable to FSQ across all metrics, particularly at higher reconstruction levels.
Analysis of GSQ’s Effectiveness Through Latent Space Geometry
Our information-theoretic analysis suggests that if GSQ preserves more structural information, this should be reflected in the geometric organization of the latent space [saxe2019information, goldfeld2020information]. Specifically, a quantization method that maintains higher mutual information and lower distortion should produce more coherent and well-separated clusters when visualized. To test this hypothesis, we employed t-SNE visualization [van2008visualizing], which maps high-dimensional data to two dimensions while preserving local neighborhood relationships, making it ideal for assessing how well structural information is maintained after quantization. Figure 5 shows striking differences between standard FSQ (Figure 5(a)) and our GSQ approach (Figure 5(b)). GSQ produces significantly more coherent clusters with clearer boundaries, indicating that GSQ preserves class structure and separability in the latent space. These results directly validate our hypothesis: GSQ’s decomposition strategy prevents information mixing across feature groups, enabling it to preserves more meaningful information within severe token constraints. The improved geometric organization of GSQ’s latent space explains its superior reconstruction quality across all our evaluation metrics.
VI-C Generalization to Out-of-Domain and Code-Switched Speech
To evaluate the generalization capability of our model beyond English and the LibriSpeech training domain, we conduct a zero-shot inference experiment on the SEAME dataset [lyu2010seame], a 200-hour spontaneous Mandarin-English code-switching corpus widely used in multilingual speech research. SEAME poses three primary challenges: (i) Conversational spontaneity: frequent disfluencies (hesitations, overlaps, false starts) and colloquial prosody; (ii) Telephony distortions and environmental noise: low-bitrate channel artifacts and background sounds typical of mobile or landline settings; (iii) Cross-lingual code-switching: rapid alternation between Mandarin and English within utterances.
Without any fine-tuning, we directly apply our Distinctive Codec model trained on LibriSpeech to a subset of 50 utterances randomly selected from SEAME. The evaluation focuses on PESQ, which serves as the primary perceptual quality metric and correlates strongly with both intelligibility and signal fidelity.
Under a low token rate setting (9.5 Hz), our model achieves a PESQ score of 1.4214, significantly outperforming SpeechTokenizer’s PESQ score of 1.0695 under the same configuration. This substantial improvement demonstrates that the proposed distinctive feature-based tokenization not only preserves critical acoustic cues in the English domain but also exhibits robust transferability to unseen languages and mixed-lingual acoustic conditions.
VI-D Comparison with WaveTokenizer
TABLE V: Performance comparison between WaveTokenizer and our Distinctive Codec at similar token rates. Lower values are better for MEL Error, STFT, and WER; higher values are better for PESQ and STOI.
Model Token Rate Metrics
[…]
WaveTokenizer 10 Hz 0.3139 1.0459 1.8333 0.7449 0.5535
Distinctive Codec (GSQ) 9.5 Hz 0.2006 0.6021 2.1901 0.8114 0.3061
To further validate the effectiveness of our distinctive feature-based approach, we conducted additional experiments comparing our method with WaveTokenizer [ji2024wavtokenizer], another state-of-the-art neural speech codec. While our main results (Table II) demonstrate competitive performance against SpeechTokenizer, this additional comparison provides broader context for our approach within the current landscape of speech tokenization methods.
For a fair comparison, we trained WaveTokenizer on the LibriSpeech dataset with a token rate of 10 Hz, matching the operating conditions of our Distinctive Codec (9.5 Hz). The results are presented in Table V.
VI-E Investigate the impact of semantic distillation